1) The following table shows the frequency of outcomes when two distinguishable coins were tossed 5,000 times and the uppermost faces were observed.

Outcome HH HT TH TT
Frequency 1,100 950 1,300 1,650

Determine the relative frequency distribution:

HH-

HT-

TH-

TT-

2) The following chart shows the results of a survey of the status of subprime home mortgages in a certain state in November 2008.

Mortgage Status Current Past Due In Foreclosure Repossessed
Frequency 110 65 60 15

(The four categories are mutually exclusive; for instance, "Past Due" refers to a mortgage whose payment status is past due but is not in foreclosure, and "In Foreclosure" refers to a mortgage that is in the process of being foreclosed but not yet repossessed. Enter your answers to two decimal places.)

(a) Find the relative frequency distribution for the experiment of randomly selecting a subprime mortgage in the state and determining its status.

Current     _______
Past Due     _______
In Foreclosure     _______
Repossessed     _______


(b) What is the relative frequency that a randomly selected subprime mortgage in the state was neither in foreclosure nor repossessed? _________

3) The following table shows crashworthiness ratings for 16 cars. (3 = Good, 2 = Acceptable, 1 = Marginal, 0 = Poor)

Frontal Crash Test Rating 3 2 1 0
Frequency 1 12 1 2

(a) Find the relative frequency distribution for the experiment of choosing a car at random and determining its frontal crash rating. (Enter your answers to four decimal places.)

Good (3)     _______
Acceptable (2)     _______
Marginal (1)     _______
Poor (0)     _______


(b) What is the relative frequency that a randomly selected car will have a crash test rating of "Marginal" or worse? (Enter your answer to four decimal places.) ______
Answer #1

1) RELATIVE FREQUENCY:

Outcome Relative frequency
HH 1100/5000 = 0.22
HT 950/5000 = 0.19
TH 1300/5000 = 0.26
TT 1650/5000 = 0.33

----------------------------

2) a) Relative frequency distribution:

Outcome Relative frequency
Current     110/250 = 0.44
Past Due     65/250 = 0.26
In Foreclosure     60/250 = 0.24
Repossessed     15/250 = 0.06

b) Relative frequency that a randomly selected subprime mortgage in the state was neither in foreclosure nor repossessed = 1 - [0.24+0.06] = 0.70

----------------------------

3) a) Relative frequency distribution:

Outcome Relative frequency
Good (3)     1/16 = 0.0625
Acceptable (2)     12/16 = 0.7500
Marginal (1)     1/16 = 0.0625
Poor (0)     2/16 = 0.1250

b) Relative frequency that a randomly selected car will have a crash test rating of "Marginal" or worse = 0.0625 + 0.1250 = 0.1875